OpenStudy (anonymous):

A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age?

OpenStudy (anonymous):

0.2x=72 x=72/02.=360 360(0.8)=288 people 30-40yrs old

OpenStudy (anonymous):

x=72/0.2*

OpenStudy (anonymous):

?

OpenStudy (anonymous):

112 288 360 432

OpenStudy (anonymous):

just made a typo up above, the answer is still the same

OpenStudy (anonymous):

288

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw